For product B, a firm has an annual holding cost that is 30% of the item price, an ordering cost of $30 per order, and annual demand of 240,000 units. This product has the following discount price ranges.

Order size

Price Per Unit

1-2499

$3.30

2500-3499

$3.20

3500-4999

$3.10

5000 or more

$3.00

-Determine the most cost-effective ordering quantity

-What is the total cost for the order quantity determined in 1)

-Describe inventory policy
Answer #1

EOQ with highest item price

EOQ = [ 2 x demand x cost od ordering / holding cost of one unit per year]1/2

= [ 2 x 240000x30/3.3x0.3]1/2 = 3814 units

This quantity falls into bracket 3, which makes the EOQ
= [ 2 x 240000x30/3.1x0.3]1/2 =3935 units

Cost with 3935 units

= cost of purchase+ cost of ordering + cost of holding

= 240000x3.1+(240000/3935) x30 + (3935/2)x 3.1x0.3

= 744000+1830+1830 =747660.

Cost with 5000 units quantity ( next bracket)

= 3x240000+30x (240000/5000)+ (5000)/2 x 3x0.3 = 723690

Ordering 5000 units will be the most economical policy for the company due to least total cost of 723690

The order size = 5000

order number ( per year) =240000/5000 =48